##############################################################
#
#  A simple example of a Maple 'source' file.  Such a file can
#  be easily created and maintained using your favorite text
#  editor and can contain arbitrary Maple commands.
#  I find this mechanism particularly useful for developing
#  and maintaining Maple procedures.
#
#  The file can most easily be read into a Maple session by
#  first 'cd'-ing to the directory which contains this file:
#
#  % cd /Public/Members/matt/Src/maple/examples
#
#  starting up maple or xmaple
#
#  % maple
#
#  or
#
#  % xmaple &
#
#  then typing
#
#  > read example;
#
#  If maple (or xmaple) isn't running in the directory
#  containing this file, then you must use an absolute
#  pathname and be sure to enclose the name in backquotes
#  or double-quotes.  (This also applies to filenames
#  containing a '.', which is why I tend to use simple
#  names (no extensions) for files containing Maple source.)
#
#  > read `/Public/Members/matt/Src/maple/examples/example`;
#

#  Recall that use of the colon (:) as terminator rather
#  than semi-colon (;) inhibits echoing of results.
#
##############################################################
aa := 23 / 155;
myadd := proc(x::numeric, y::numeric)
        x + y;
end;